OpenStudy (anonymous):

I could really use some help please. Simplify 6P2?

OpenStudy (mathstudent55):

\(_nP_r = \dfrac{n!}{(n - r)!}\)

OpenStudy (mathstudent55):

The first step is copy the formula. Then copy the formula again replacing n and r with your values. Show your work. If you need help evaluating, just ask.

OpenStudy (anonymous):

I agree with mathstudent55

OpenStudy (anonymous):

Thank you =D This helps

OpenStudy (mathstudent55):

@DasiabemynameBVB Cna you finish it and get a value?

OpenStudy (anonymous):

I'm sorry I'm actually very confused

OpenStudy (mathstudent55):

Step 1. Copy the formula: \(_nP_r = \dfrac{n!}{(n - r)!}\) Step 2. Copy the formula again, but substitute your values of n and r: \(_6P_2 = \dfrac{6!}{(6 - 2)!}\) Step 3. Evaluate the expression. \(_6P_2 = \dfrac{6 \times 5 \times \cancel{4!}}{\cancel{4!}} = 6 \times 5 = 30\)